OpenStudy (anonymous):

Simplify the square root of 9+ the square root of 16

OpenStudy (anonymous):

3+4=7,try these urself please :)

OpenStudy (anonymous):

root9 =3 root16 = 4 so 3+4=7

OpenStudy (anonymous):

\[\sqrt{9+16}\]

OpenStudy (anonymous):

this is 5

OpenStudy (anonymous):

i.e. for the above ans is 5 because 9+16 = 25 and root25 is 5